# ProS: Prompting-to-simulate Generalized knowledge for Universal Cross-Domain Retrieval

### 19 Dec 2023 v1

@@ Abstract
The goal of Universal Cross-Domain Retrieval (UCDR) is to achieve robust performance in generalized test scenarios, wherein data may belong to strictly unknown domains and categories during training. Recently, pre-trained models with prompt tuning have shown strong generalization capabilities and attained noteworthy achievements in various downstream tasks, such as few-shot learning and video-text retrieval. However, applying them directly to UCDR may not sufficiently to handle both domain shift (i.e., adapting to unfamiliar domains) and semantic shift (i.e., transferring to unknown categories). To this end, we propose Prompting-to-Simulate (ProS), the first method to apply prompt tuning for UCDR. ProS employs a two-step process to simulate Content-aware Dynamic Prompts (CaDP) which can impact models to produce generalized features for UCDR. Concretely, in Prompt Units Learning stage, we introduce two Prompt Units to individually capture domain and semantic knowledge in a mask-and-align way. Then, in Context-aware Simulator Learning stage, we train a Content-aware Prompt Simulator under a simulated test scenarios to produce the corresponding CaDP. Extensive experiments conducted on three benchmark datasets show that our method achieves new state-of-the-art performance without bringing excessive parameters. Our method is publicly available at https://anonymous.4open.science/r/ProS
@@ 1 Introduction
Cross-Domain Retrieval (CDR) [43, 11] is an important task in Information Retrieval (IR) systems that utilizes data from one domain (e.g. Infograph, Sketch) as a query to search semantically similar examples in another domain (e.g. natural image). Despite achieving promising results, existing methods [34, 11, 2, 4] heavily follow a close-set learning setting, where domains and categories in both training and testing data are pre-defined. Such evaluation limiting the model to the training samples and unable to tackle the open-set applications such as e-commerce search [6, 10, 25] and recommendation [18, 10], where searched product not only comes from various new domains (e.g. styles) but is also accompanied by the emergence of new categories. Towards this goal, one emerging research direction is proposed, namely Universal Cross-Domain Retrieval (UCDR) [29], shown in Fig. 1(a), for performing well under generalized test scenarios, where test data may belong to strictly unknown domains and categories.
The core challenge for UCDR is how to empower models with the capabilities of both domain shift (i.e. adapting to unfamiliar domains) and semantic shift (i.e. transferring to unknown categories). The mainstream approaches [29, 37] focus on fine-tuning visual models, e.g. ResNet [14] and ViT [8], with the guidance of expert knowledge from semantic information or pre-trained models. Compared with ResNet, ViT has stronger representational capability and contains more knowledge gained from larger pre-training data, which facilitates the generalization of the model for unknown images. This raises a intuitive question: does the extensive general knowledge inherent in large-scale pre-trained models can enhance it’s ability to tackle UCDR task where priori information about test samples are strict unknown?
With this question in mind, we empirically conduct a confirmatory experiment on the widely-recognized large-scale pre-trained model, CLIP [31] under UCDR. Specifically, we full fine-tune the CLIP and compare its performance with ViT and ResNet, as shown in Fig. 1 (b). From the figure, we can see that CLIP yields tremendous gains, reaching a 22.27% improvement in mAP@200 compared with ViT. This result proves the effectiveness of common knowledge inherent in large-scale pre-trained CLIP for UCDR task.
@@
Towards above question, we propose a novel prompt tuning method in a simulated way, named Prompting-to-Simulate (ProS), to effectively leverage prompt tuning to mine the generalized knowledge from CLIP for UCDR. Specifically, ProS introduces a two-stage learning to simulate two content-aware dynamic prompts (CaDP) which reflect input sample’s domain and category, respectively. In Prompt Unit Learning (PUL), two groups of distinct learnable prompts are established: domain prompt units and semantic prompt units. These two units are employed to extract domain knowledge and semantic knowledge from source data in a mask-and-align strategy. Then, in Context-aware Simulator learning (CSL) , we train a Content-aware Prompt Simulator (CaPS) under simulated test scenarios to dynamically generate content-aware dynamic prompts (CaDP) based on the learned prompt units. With the CaDP, we can impact CLIP’s knowledge to gain more generalizable representation and achieve better performance compared with existing prompt tuning methods in UCDR, as shown in Fig. 1 (b). In addition, our method uses considerably fewer learnable parameters compared to full fine-tuning, accounting for only 12.84% of the model parameters, and does not introduce excessive learnable parameters compared to prompt tuning methods.
@@
  • We are the first to investigate how to adapt CLIP with prompt for UCDR.
@@ 2.1 Universal Cross-Domain Retrieval
UCDR requires solving domain and category generalization simultaneously which can be roughly regarded as a combination of Domain Generalization (DG) [24, 38] and Zero-Shot Learning (ZSL) [39, 13]. In DG, some approaches synthesize unseen samples using data augmentation strategy [28] or generative adversarial networks (GANs) [40] to handle the domain shift. In ZSL, some researchers leverage the auxiliary data, including human-annotated attribute information  [21, 20], text description[32] or knowledge graph [22] to learn relationship between seen categories and pre-define unseen categories names. It should be clarified that UCDR does not require pre-define unseen class names. Inspired by above approaches, SnMpNet [29] trains the model with the mix-up data augmentation strategy and align features with semantic information. Additionally, SASA[37] deploys a ViT-based model and train categorical prototypes to handle the semantic shift. Different from these works, we make the first attempt to apply a large-scale pre-trained model, i.e. CLIP, to the UCDR task, and proposed a new prompt tuning method to improve the generalization ability of the CLIP.
@@ 2.2 Vision-Language Pre-training Models
Motivated by the success in self-supervised learning, Vision-Language Pre-training has emerged as a prominent topic. CLIP [31] initially demonstrated that equipped with large-scale image-text pairs, contrastive learning framework can achieve a performance comparable to fully supervised baselines. Subsequently, ALIGN [16] scale the training dataset to billions and gain a better vision-language representation. Another line of works [12, 5] leverage pre-trained object detection models like Faster-RCNN [33] to extract image regional features offline for training multi-modal transformers. These models show impressive performance on various vision-language tasks, as they are capable of processing both single-modal and multi-modal inputs. In this paper, we make the first attempt to leverage the powerful CLIP for the UCDR task.
@@ 2.3 Prompt Tuning
Prompt Tuning [27], as a new paradigm, initially surface in NLP for adapting pre-trained language models (PLMs) [7, 3] to downstream tasks. These prompts can either be handcrafted for specific tasks [36] or learned automatically via gradients which termed as “Prompt Learning” [23]. Driven by the success of prompts paradigm in NLP, various studies like CoOP [41] and CoCoOp [42] apply text prompting in multi-modal scenarios, while VPT [17] and VP [1] introduces these paradigm into vision transformer. Despite the pioneering successes of VP and VPT in visual models, we find that they are static after training which is not flexible to handle the generalized scenarios. In contrast, our method propose a Prompt-to-Simulate method to well address UCDR task.
@@ 3 Method
Our method makes the first attempt to apply CLIP for Universal Cross-Domain Retrieval (UCDR) with a novel prompt learning method named ProS, as shown in Fig. 2. We begin with a brief overview of some important preliminaries about UCDR, CLIP and prompt tuning paradigm (Sec. 3.1). Then, a detailed description of our proposed Prompting-to-Simulate (Pros) is provided (Sec. 3.2), including two training stages: Prompt Unit Learning (PUL) and Context-aware Simulator Learning (CSL). Finally, we will introduce how to perform retrieval using our proposed method (Sec. 3.3).
Figure 2: Overview of our proposed ProS. In Prompt Units Learning Stage, we capture knowledge from source data into domain prompts units $DP$ and semantic prompts units $SP$ by masking irrelevance prompts. In Context-aware Prompt Simulation Stage, we train a Context-aware Prompt Simulator (CaPS) with a mask operation to dynamically convey prompt templates $PT$ to two Content-aware Dynamic Prompts (CaDP) to simulate unknown domains and categories. In retrieval phase, we employ CaPS to produce CaDP which impact the CLIP image encoder to convert unseen samples into suitable embeddings for retrieval. The gray parts indicate masked prompts.
@@ 3.1 Preliminary
Problem Formulation. Universal Cross-Domain Retrieval (UCDR) [29] aims to accurately retrieve images with queries from unseen domains and classes. From a general perspective, we assume that training dataset is composed of $K$-different ($K\geq 2$) source domains (Real, Sketch, Quickdraw, etc.) which can be formulated as $\mathcal{D}_{train}=\bigcup_{d\in\left\{1,\ldots,K\right\}}\left\{\boldsymbol{x}_{i}^{d},y^{tr}_{i}\right\}_{i=1}^{N_{d}}$. The $d^{th}$ domain has $N_{d}$ images, where each image $\boldsymbol{x}_{i}^{d}$ is belong to class $y^{tr}_{i}\in\mathcal{C}_{train}$. In the testing phase, the gallery set and query set are needed to enable image retrieval. Specifically, query set $\mathcal{D}_{query}=\left\{\boldsymbol{x}_{i}^{q},y^{te}_{i}\right\}_{i=1}^{N_{q}}$ consists of $N_{q}$ samples belong to unseen class $y_{i}^{\text{te}}\in\mathcal{C}_{\text{test}}$ from the query domain $q$, while gallery set consists of real images. To simulate a more realistic retrieval scenario, the gallery set has two settings: 1) all samples belong to unseen class, termed Unseen Gallery; (2) samples belong to both seen class and unseen class, termed Mixed Gallery.
The UCDR protocol can be separated into two sub-tasks, namely: (a) $\mathrm{U}^{c}\mathrm{CDR}$, where the query domain $q\in\{1,...,K\}$ but the classes are held out, i.e. $\mathcal{C_{\text{train}}}\cap\mathcal{C_{\text{test}}}\equiv\phi$. (b) $\mathrm{U}^{d}\mathrm{CDR}$, where the query domain $q\notin\{1,...,K\}$ but the classes are seen, i.e. $\mathcal{C_{\text{train}}}\equiv\mathcal{C_{\text{test}}}$. The goal of UCDR is to learn a mapping function that can place query and gallery images into a latent domain-independent subspace $\Phi$ , where images of same label but different domains are clustered together. The most challenging part is images from unseen domains and unknown classes also need to be mapped correctly in $\Phi$.
@@
where $\otimes$ is cosine similarity. Benefiting from large-scale pre-training and alignment of images and text, CLIP is robust for various visual appearance changes [15], which is suitable for UCDR task. However, fine-tuning CLIP will distort pre-trained features and lead to poor out-of-distribution performance [19]. Therefore, how to effectively leverage this powerful foundation model for UCDR becomes an important problem.
Prompt Tuning. Compared with full fine-tuning, prompt tuning prevents pre-trained feature distortion by introducing task-specific trainable parameters termed prompts into the input while keeping pre-trained model frozen. Specifically, [41] incorporate learnable text prompts $\mathbf{P_{t}}=\left\{p_{t}^{i}\in\mathbb{R}^{\ell}\right\}_{i=1}^{N_{t}}$ into the text template as the input of CLIP text encoder, where $N_{t}$ and $\mathbb{R}^{\ell}$ indicate text prompts length and dimensions, respectively. Then the text template can be represented as “$\mathbf{P_{t}}$ a photo of a [CLASS]”. Furthermore, [17] proposed visual prompts for ViT which can be formulated as $\mathbf{P_{v}}=\left\{p_{v}^{i}\in\mathbb{R}^{\ell}\right\}_{i=1}^{N_{v}}$, where $N_{v}$ means the number of visual prompts and $p_{v}$ is a learnable parameter with $\ell$ dimension. Then the visual prompts are inserted into the first Transformer layer, named VPT-Shallow in [17] :
@@
where $\mathbf{Z}_{i}\in\mathbb{R}^{N_{v}\times\ell}$ represents the prompt embeddings computed by the $i^{th}$ Transformer layer in the CLIP image encoder, $\mathbf{E}_{0}$ means the image patch embeddings extracted from a CNN feature map [8]. In addition, $\mathbf{x_{0}}$ is CLS token, which produces $\mathbf{y}$, and served as the final output of image representation.
@@ 3.2 Prompt-to-Simulate
where $\odot$ discards prompts with corresponding mask of 00 preserves others of $1$. Correspondingly, image feature is obtained in the same way as Eq. 2. Since text feature can be considered as the class prototype, we introduce the learnable text prompts $\mathbf{P_{t}}$ as we discussed in Sec. 3.1 to achieve more flexible training. Additionally, we design a new text template: “a photo of $[CLASS]^{i}$ from $\mathbf{P_{t}}$ domain.”, to better suit the UCDR task. Given the prompted input $\mathbf{x_{p}}$ and the class caption $\left\{\boldsymbol{t}_{i}\right\}_{i=1}^{|C_{train}|}$ generated by above text template, the training loss can be formulated as:
@@
Context-aware Simulator Learning (CSL). In CSL, our goal is to train a module termed Content-aware Prompt Simulator (CaPS) to generate two Content-aware Dynamic Prompts (CaDP) similar to Prompt Units which can well impact CLIP to extract generalized features of unseen samples. Towards this goal, we concatenate two sets of Prompt Units with image patches as CaPS’s input, which can be formulate as $[PT_{d},PT_{s},DP,SP,\mathbf{E_{0}}]$, where $PT_{d}\in\mathbb{R}^{\ell},\;PT_{s}\in\mathbb{R}^{\ell}$ are domain and semantic prompt template for CaDP generation and $\mathbf{E_{0}}$ are unseen sample’s embedding.
Inspired by the ieda of meta-learning, we train CaPS under a simulated test scenario by a relevant mask operation. Specifically, we mask relevant domain and semantic prompt units w.r.t. input samples by $\mathbf{M_{rd}}=\left\{\mathbf{m_{rd}}^{i}\in[0,1]\right\}^{K}_{i=1}$ and $\mathbf{M_{rs}}=\left\{\mathbf{m_{rs}}^{i}\in[0,1]\right\}^{\left|\mathcal{C_{\text{train}}}\right|}_{i=1}$, where $\mathbf{m_{rd}^{d}}=0$, $\mathbf{m_{rs}^{c}}=0$ and others are $1$ if the input sample is from domain $\mathbf{d}$ and class $\mathbf{c}$. Then we can hide the prompt units corresponding to input samples by multiplying them with $\mathbf{M_{rs}}$ and $\mathbf{M_{rs}}$ mask matrices. The whole process can be formulated as:
@@
Similarly, we use the same objective as Eq. 6 to train the newly introduced $PT_{d},PT_{s}$ and $\mathcal{M}$ while remaining part would be fixed in this stage.

### 29 Feb 2024 v3

@@ Abstract
The goal of Universal Cross-Domain Retrieval (UCDR) is to achieve robust performance in generalized test scenarios, wherein data may belong to strictly unknown domains and categories during training. Recently, pre-trained models with prompt tuning have shown strong generalization capabilities and attained noteworthy achievements in various downstream tasks, such as few-shot learning and video-text retrieval. However, applying them directly to UCDR may not be sufficient to handle both domain shift (i.e., adapting to unfamiliar domains) and semantic shift (i.e., transferring to unknown categories). To this end, we propose Prompting-to-Simulate (ProS), the first method to apply prompt tuning for UCDR. ProS employs a two-step process to simulate Content-aware Dynamic Prompts (CaDP) which can impact models to produce generalized features for UCDR. Concretely, in Prompt Units Learning stage, we introduce two Prompt Units to individually capture domain and semantic knowledge in a mask-and-align way. Then, in Context-aware Simulator Learning stage, we train a Content-aware Prompt Simulator under a simulated test scenario to produce the corresponding CaDP. Extensive experiments conducted on three benchmark datasets show that our method achieves new state-of-the-art performance without bringing excessive parameters. Our method is publicly available at https://github.com/fangkaipeng/ProS
@@ 1 Introduction
Cross-Domain Retrieval (CDR) [43, 11] is an important task in Information Retrieval (IR) systems that utilizes data from one domain (e.g. Infograph, Sketch) as a query to search semantically similar examples in another domain (e.g. natural image). Despite achieving promising results, existing methods [34, 11, 2, 4] heavily follow a close-set learning setting, where domains and categories in both training and testing data are pre-defined. Such evaluation limits the model to the training samples and unable to tackle the open-set applications such as e-commerce search [6, 10, 25] and recommendation [18, 10], where a searched product not only comes from various new domains (e.g. styles) but is also accompanied by the emergence of new categories. Towards this goal, one emerging research direction is proposed, namely Universal Cross-Domain Retrieval (UCDR) [29], shown in Fig. 1(a), for performing well under generalized test scenarios, where test data may belong to strictly unknown domains and categories.
The core challenge for UCDR is how to empower models with the capabilities of both domain shift (i.e. adapting to unfamiliar domains) and semantic shift (i.e. transferring to unknown categories). The mainstream approaches [29, 37] focus on fine-tuning visual models, e.g. ResNet [14] and ViT [8], with the guidance of expert knowledge from semantic information or pre-trained models. Compared with ResNet, ViT has stronger representational capability and contains more knowledge gained from larger pre-training data, which facilitates the generalization of the model for unknown images. This raises an intuitive question: can the extensive general knowledge inherent in large-scale pre-trained models enhance its ability to tackle the UCDR task where prior information about test samples is strictly unknown?
With this question in mind, we empirically conduct a confirmatory experiment on the widely-recognized large-scale pre-trained model, CLIP [31] under UCDR. Specifically, we fully fine-tune the CLIP and compare its performance with ViT and ResNet, as shown in Fig. 1 (b). From the figure, we can see that CLIP yields tremendous gains, reaching a 22.27% improvement in mAP@200 compared with ViT. This result proves the effectiveness of common knowledge inherent in large-scale pre-trained CLIP for UCDR task.
@@
Towards above question, we propose a novel prompt tuning method in a simulated way, named Prompting-to-Simulate (ProS), to effectively leverage prompt tuning to mine the generalized knowledge from CLIP for UCDR. Specifically, ProS introduces a two-stage learning to simulate two content-aware dynamic prompts (CaDP) which reflect the input sample’s domain and category, respectively. In Prompt Unit Learning (PUL), two groups of distinct learnable prompts are established: domain prompt units and semantic prompt units. These two units are employed to extract domain knowledge and semantic knowledge from source data in a mask-and-align strategy. Then, in Context-aware Simulator learning (CSL), we train a Content-aware Prompt Simulator (CaPS) under simulated test scenarios to dynamically generate content-aware dynamic prompts (CaDP) based on the learned prompt units. With the CaDP, we can impact CLIP’s knowledge to gain a more generalizable representation and achieve better performance compared with existing prompt tuning methods in UCDR, as shown in Fig. 1 (b). In addition, our method uses considerably fewer learnable parameters compared to full fine-tuning, accounting for only 12.84% of the model parameters, and does not introduce excessive learnable parameters compared to prompt tuning methods.
@@
  • We are the first to investigate how to adapt CLIP with prompts for UCDR.
@@ 2.1 Universal Cross-Domain Retrieval
UCDR requires solving domain and category generalization simultaneously which can be roughly regarded as a combination of Domain Generalization (DG) [24, 38] and Zero-Shot Learning (ZSL) [39, 13]. In DG, some approaches synthesize unseen samples using data augmentation strategy [28] or generative adversarial networks (GANs) [40] to handle the domain shift. In ZSL, some researchers leverage the auxiliary data, including human-annotated attribute information  [21, 20], text description[32] or knowledge graph [22] to learn the relationship between seen categories and pre-define unseen categories names. It should be clarified that UCDR does not require pre-defined unseen class names. Inspired by above approaches, SnMpNet [29] trains the model with the mix-up data augmentation strategy and align features with semantic information. Additionally, SASA[37] deploys a ViT-based model and trains categorical prototypes to handle the semantic shift. Different from these works, we make the first attempt to apply a large-scale pre-trained model, i.e. CLIP, to the UCDR task, and propose a new prompt tuning method to improve the generalization ability of the CLIP.
@@ 2.2 Vision-Language Pre-training Models
Motivated by the success of self-supervised learning, Vision-Language Pre-training has emerged as a prominent topic. CLIP [31] initially demonstrated that equipped with large-scale image-text pairs, a contrastive learning framework can achieve a performance comparable to fully supervised baselines. Subsequently, ALIGN [16] scales the training dataset to billions and gains a better vision-language representation. Another line of work [12, 5] leverage pre-trained object detection models like Faster-RCNN [33] to extract image regional features offline for training multi-modal transformers. These models show impressive performance on various vision-language tasks, as they are capable of processing both single-modal and multi-modal inputs. In this paper, we make the first attempt to leverage the powerful CLIP for the UCDR task.
@@ 2.3 Prompt Tuning
Prompt Tuning [27], as a new paradigm, initially surfaces in NLP for adapting pre-trained language models (PLMs) [7, 3] to downstream tasks. These prompts can either be handcrafted for specific tasks [36] or learned automatically via gradients which are termed as “Prompt Learning” [23]. Driven by the success of prompt paradigm in NLP, various studies like CoOP [41] and CoCoOp [42] apply text prompting in multi-modal scenarios, while VPT [17] and VP [1] introduce this paradigm into vision transformer. Despite the pioneering successes of VP and VPT in visual models, we find that they are static after training which is not flexible to handle the generalized scenarios. In contrast, our method proposes a Prompt-to-Simulate method to well address UCDR task.
@@ 3 Method
Our method makes the first attempt to apply CLIP for Universal Cross-Domain Retrieval (UCDR) with a novel prompt learning method named ProS, as shown in Fig. 2. We begin with a brief overview of some important preliminaries about UCDR, CLIP, and prompt tuning paradigm (Sec. 3.1). Then, a detailed description of our proposed Prompting-to-Simulate (Pros) is provided (Sec. 3.2), including two training stages: Prompt Unit Learning (PUL) and Context-aware Simulator Learning (CSL). Finally, we will introduce how to perform retrieval using our proposed method (Sec. 3.3).
Figure 2: Overview of our proposed ProS. In Prompt Units Learning Stage, we capture knowledge from source data into domain prompts units $DP$ and semantic prompts units $SP$ by masking irrelevance prompts. In the Context-aware Prompt Simulation Stage, we train a Context-aware Prompt Simulator (CaPS) with a mask operation to dynamically convey prompt templates $PT$ to two Content-aware Dynamic Prompts (CaDP) to simulate unknown domains and categories. In the retrieval phase, we employ CaPS to produce CaDP which impacts the CLIP image encoder to convert unseen samples into suitable embeddings for retrieval. The gray parts indicate masked prompts.
@@ 3.1 Preliminary
Problem Formulation. Universal Cross-Domain Retrieval (UCDR) [29] aims to accurately retrieve images with queries from unseen domains and classes. From a general perspective, we assume that the training dataset is composed of $K$-different ($K\geq 2$) source domains (Real, Sketch, Quickdraw, etc.) which can be formulated as $\mathcal{D}_{train}=\bigcup_{d\in\left\{1,\ldots,K\right\}}\left\{\boldsymbol{x}_{i}^{d},y^{tr}_{i}\right\}_{i=1}^{N_{d}}$. The $d^{th}$ domain has $N_{d}$ images, where each image $\boldsymbol{x}_{i}^{d}$ belongs to class $y^{tr}_{i}\in\mathcal{C}_{train}$. In the testing phase, the gallery set and query set are needed to enable image retrieval. Specifically, the query set $\mathcal{D}_{query}=\left\{\boldsymbol{x}_{i}^{q},y^{te}_{i}\right\}_{i=1}^{N_{q}}$ consists of $N_{q}$ samples belonging to unseen classes $y_{i}^{\text{te}}\in\mathcal{C}_{\text{test}}$ from the query domain $q$, while the gallery set consists of real images. To simulate a more realistic retrieval scenario, the gallery set has two settings: 1) all samples belong to the unseen class, termed Unseen Gallery; (2) samples belong to both seen class and unseen class, termed Mixed Gallery.
The UCDR protocol can be separated into two sub-tasks, namely: (a) $\mathrm{U}^{c}\mathrm{CDR}$, where the query domain $q\in\{1,...,K\}$ but the classes are held out, i.e. $\mathcal{C_{\text{train}}}\cap\mathcal{C_{\text{test}}}\equiv\phi$. (b) $\mathrm{U}^{d}\mathrm{CDR}$, where the query domain $q\notin\{1,...,K\}$ but the classes are seen, i.e. $\mathcal{C_{\text{train}}}\equiv\mathcal{C_{\text{test}}}$. The goal of UCDR is to learn a mapping function that can place query and gallery images into a latent domain-independent subspace $\Phi$ , where images of the same label but different domains are clustered together. The most challenging part is images from unseen domains and unknown classes also need to be mapped correctly in $\Phi$.
@@
where $\otimes$ is cosine similarity. Benefiting from large-scale pre-training and alignment of images and text, CLIP is robust for various visual appearance changes [15], which is suitable for UCDR. However, fine-tuning CLIP will distort pre-trained features and lead to poor out-of-distribution performance [19]. Therefore, how to effectively leverage this powerful foundation model for UCDR becomes an important problem.
Prompt Tuning. Compared with full fine-tuning, prompt tuning prevents pre-trained feature distortion by introducing task-specific trainable parameters termed prompts into the input while keeping the pre-trained model frozen. Specifically, [41] incorporate learnable text prompts $\mathbf{P_{t}}=\left\{p_{t}^{i}\in\mathbb{R}^{\ell}\right\}_{i=1}^{N_{t}}$ into the text template as the input of CLIP text encoder, where $N_{t}$ and $\mathbb{R}^{\ell}$ indicate text prompts length and dimensions, respectively. Then the text template can be represented as “$\mathbf{P_{t}}$ a photo of a [CLASS]”. Furthermore, [17] proposed visual prompts for ViT which can be formulated as $\mathbf{P_{v}}=\left\{p_{v}^{i}\in\mathbb{R}^{\ell}\right\}_{i=1}^{N_{v}}$, where $N_{v}$ represents the number of visual prompts and $p_{v}$ is a learnable parameter with $\ell$ dimension. Then the visual prompts are inserted into the first Transformer layer, named VPT-Shallow in [17] :
@@
where $\mathbf{Z}_{i}\in\mathbb{R}^{N_{v}\times\ell}$ represents the prompt embeddings computed by the $i^{th}$ Transformer layer in the CLIP image encoder, $\mathbf{E}_{0}$ means the image patch embeddings extracted from a CNN feature map [8]. In addition, $\mathbf{x_{0}}$ is CLS token, which produces $\mathbf{y}$ and serves as the final output of image representation.
@@ 3.2 Prompt-to-Simulate
where $\odot$ discards prompts with corresponding mask of 00 preserves others of $1$. Correspondingly, the image feature is obtained in the same way as Eq. 2. Since the text feature can be considered as the class prototype, we introduce the learnable text prompts $\mathbf{P_{t}}$ as we discussed in Sec. 3.1 to achieve more flexible training. Additionally, we design a new text template: “a photo of $[CLASS]^{i}$ from $\mathbf{P_{t}}$ domain.”, to better suit the UCDR task. Given the prompted input $\mathbf{x_{p}}$ and the class caption $\left\{\boldsymbol{t}_{i}\right\}_{i=1}^{|C_{train}|}$ generated by above text template, the training loss can be formulated as:
@@
Context-aware Simulator Learning (CSL). In CSL, our goal is to train a module termed Content-aware Prompt Simulator (CaPS) to generate two Content-aware Dynamic Prompts (CaDP) similar to Prompt Units which can well impact CLIP to extract generalized features of unseen samples. Towards this goal, we concatenate two sets of Prompt Units with image patches as CaPS’s input, which can be formulated as $[PT_{d},PT_{s},DP,SP,\mathbf{E_{0}}]$, where $PT_{d}\in\mathbb{R}^{\ell},\;PT_{s}\in\mathbb{R}^{\ell}$ are domain and semantic prompt template for CaDP generation and $\mathbf{E_{0}}$ are unseen sample’s embedding.
Inspired by the idea of meta-learning, we train CaPS under a simulated test scenario by a relevant mask operation. Specifically, we mask relevant domain and semantic prompt units w.r.t. input samples by $\mathbf{M_{rd}}=\left\{\mathbf{m_{rd}}^{i}\in[0,1]\right\}^{K}_{i=1}$ and $\mathbf{M_{rs}}=\left\{\mathbf{m_{rs}}^{i}\in[0,1]\right\}^{\left|\mathcal{C_{\text{train}}}\right|}_{i=1}$, where $\mathbf{m_{rd}^{d}}=0$, $\mathbf{m_{rs}^{c}}=0$ and others are $1$ if the input sample is from domain $\mathbf{d}$ and class $\mathbf{c}$. Then we can hide the prompt units corresponding to input samples by multiplying them with $\mathbf{M_{rd}}$ and $\mathbf{M_{rs}}$ mask matrices. The whole process can be formulated as:
@@
Similarly, we use the same objective as Eq. 6 to train the newly introduced $PT_{d},PT_{s}$ and $\mathcal{M}$ while the remaining part would be fixed in this stage.